OpenStudy (anonymous):

what is the slope of the line that connects the points (2,3) and (-4,8)

OpenStudy (anonymous):

use the equation y2-y1/x2-x1

OpenStudy (anonymous):

Slope is given by \[y _{2} - y _{1} \div x _{2} - x _{1}\] So the slope is -5/6

OpenStudy (anonymous):

This is the equation by which you can find the slope  Make data that will help you alot :) The first bracket will be x1 and y1 (2,3) The second bracket is x2 and y2 (-4,8) Data: x1=2 x2=-4 y1=3 y2=8 y2-y1/x2-x1 = 8-3/-4-2 = 5/-6 So, the slope is -5/6